MILLIMETER-WAVE SPECTRUM OF FeNO(X^{2}Δ_{i}) IN THE GROUND AND VIBRATIONALLY EXCITED STATES

MOTOKI NAKASHIMA, MASATO HAYASHI, KENSUKE HARADA and \underline{KEIICHI TANAKA}, Department of Chemistry, Faculty of Science, Kyushu University 33, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

Rotational spectrum of the FeNO radical generated by the ultraviolet photolysis of Fe(CO)_{2}(NO)_{2} was measured in the millimeter-wave region. The measurements were performed in the supersonic jet expansion with a millimeter-wave multi-reflection cell. Three rotational transitions (J = 9.5 − 8.5 ∼ 11.5 − 10.5) in the Ω = 5/2 spin substate of the X^{2}Δ_{i} ground vibronic state were measured in the frequency region of 87−106 GHz. The rotational lines were split into 2 components (ΔF = 0, +1) due to the hyperfine interaction of the N(I = 1) nucleus. In the upper spin substate Ω = 3/2 (A_{SO} ≈ −417 cm^{−1}) of the electronic ground state, seven rotational transitions (J = 28.5 − 27.5 ∼ 34.5 − 33.5) were measured with a conventional absorption cell (2.7 m in length) in the room temperature. Moreover, six rotational transitions (J = 28.5 − 27.5 ∼ 33.5 − 32.5) in the ν_{2} vibrationally excited state (for both P = 3/2 and 7/2 components) and the high-J lines of Ω = 5/2 spin substate of the ground state were also observed with the conventional absorption cell. Rotational line intensity of the Ω = 3/2 substate was about one tenth of that for the Ω = 5/2 substate because of the large spin-orbit interaction constant A_{SO}. Molecular constants, including the rotational constant B, centrifugal distortion constant D, hyperfine constant a + b_{F} /4 + c/6, and vibration rotation constant α_{2}, were determined by a least squares fitting of the observed spectrum. The electronic ground state of FeNO was confirmed to be X^{2}Δ_{i} as in the case of CoCO and the unpaired electron is localized almost in the 3d orbital of Fe. The hyperfine constant of FeNO, a + b_{F} /4 + c/6 = −1.359(57) MHz, is much smaller than that of CoCO, 466.073(54) MHz. Rovibrational transitions were also observed by the infrared diode laser spectroscopy with the ultraviolet photolysis of Fe(CO)_{2}(NO)_{2}.^{a}

aRadicals and Ions session in the present symposium.